Question 41 ptsSkip to question text.

At constant pressure, density of gas is inversely proportional to its temperature. This is how hot air balloon is floating in air: burning fuel makes inside air temperature higher and density lower; low density air balloon float on high density regular air. What is the density of air in g/L at 76 degree C if the atmospheric pressure is 1.0 atm? The average molar mass of air is about 29 g/mol. Keep two decimal places for your answer.
Answer #1

MW air = 29

then

Dair1 = ?

when

T1 = 76°C = 349 K

P1 = 1 atm

PV = nRT

mass = mol/MW

D = mass/V

PV = nRT

PV = mass/MW*RT

mass/V = P*MW/(RT)

D = P*MW/(RT)

D = (1)(29)/(0.082*349) = 1.013348 g/L

D = 1.013348 g/L